An airplane travels 260 miles per hour into a 50 mile per hour head wind. On the return trip the wind has slowed to 25 miles per hour and the plane averages 335 miles per hour. How fast would the plane have gone with no wind on either leg?
Found 2 solutions by checkley77, ankor@dixie-net.com:
Answer by checkley77(12844)

260+50=335-25
310=310 is the speed of the plane with no wind.

Answer by ankor@dixie-net.com(22740)

An airplane travels 260 miles per hour into a 50 mile per hour head wind.
On the return trip the wind has slowed to 25 miles per hour and the plane averages 335 miles per hour.
How fast would the plane have gone with no wind on either leg?
:
Let s = speed of plane in still air
then given that:
(s-50) = 260
s = 260 + 50
s = 310 mph in still air
and return trip
(s+25) = 335
s + 25 = 335
s = 335 - 25
s = 310 mph, as expected